A home was purchased for $120,000 in 2001. After one year the home had appreciated 5%. After one more year, the home had appreciated an additional 4%. What was the value of the home after these two years, rounded to the nearest hundred dollars?
A) $124,800 B) $126,000 C) $130,800 D) $131,000
. A new car is marked down from $28,000 to $25,200. What is the discount rate?
A) 9% B) 11% C) 10% D) 15%

Answer by josmiceli(19441)

P = purchase price
a = rate of appreciation
V(1) = value after 1 year
V(2) = value after 2 years
The first year, a = 5%
P = 120000
V(1) = P + aP
V(1) = 120000 + (.05)(120000)
V(1) = 120000(1 + .05)
V(1) = 126000
V(2) = V(1) + aV(1)
V(2) = 126000 + (.04)(126000)
V(2) = 126000(1 + .04)
V(2) = 131040
the answere is (d) 131,000
discount rate = (original price) - (discounted price) / (original price)
28000 - 25200 = 2800
2800 / 28000 = .10
the answer is (c) 10%